OpenStudy (anonymous):

Find dy/dx by implicit differentiation. 4 x**2 - 5 y**2 = 16

OpenStudy (cwrw238):

what do you mean by ** ?

OpenStudy (anonymous):

4x^2-5y^2=16

OpenStudy (cwrw238):

is it a power?

OpenStudy (cwrw238):

oh ok

OpenStudy (anonymous):

i got dy/dx = 8x/5y

OpenStudy (anonymous):

4x/5y actually

OpenStudy (cwrw238):

when you differentiate implicitly (dy/dx) you treat y as a function of x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got teh answer i will bump up your score

OpenStudy (cwrw238):

8x - 5* 2y* dy*dx = 0 sd/dx = 8x/10y = 4x/5y

OpenStudy (anonymous):

ty

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

how do i give you score

OpenStudy (cwrw238):

you apply the chain rule to y^2

OpenStudy (cwrw238):

just click on best reponse

OpenStudy (anonymous):

really? i thought it was only the power rule

OpenStudy (anonymous):

wouldnt this be chain rule (4x^2)^2

OpenStudy (cwrw238):

yes - but you are taking y to be a function of x so it amounts to the same thing

OpenStudy (cwrw238):

to be honest i cant remember the power rule plz remind me